# Towards the Worst-case Robustness of Large Language Models

### v3

@@ D.4.3 Simple case study: $d$=1,2,3
This motivate us to provide the general proof in Eq. 10
@@ E.1 Reduction Lemma
Therefore, the certified bound of the uniform kernel is independent of the input length $N$ (dependent part only exists in network accuracy $p_{A}$), but only adversarial budget $d$. This greatly simplifies the derivation of value-to-weight ratio and volume. We give these results in the following Theorem 5.4. We can compute the certified robustness using the uniform kernel by plugging these results into Algorithm 1.
@@ E.2 Symmetrization Lemma
can only take $12d+$ 1 distinct values, specifically:
@@
$=\binom{d}{i}\binom{i}{i+j-d}(|\mathcal{V}|-2)^{i+j-d}\cdot\alpha^{j}\bar{\beta}^{d-j}$ by ($\text{by }\binom{A}{B}=\binom{A}{A-B}$
$=\binom{d}{i+j-d}\binom{2d-i-j}{d-j}(|\mathcal{V}|-2)^{i+j-d}\cdot\alpha^{j}\bar{\beta}^{d-j}$ by ($\text{by }\binom{A}{B}\binom{B}{C}=\binom{A}{C}\binom{A-C}{B-C}$
$=\binom{d}{i+j-d}\binom{2d-i-j}{d-i}(|\mathcal{V}|-2)^{i+j-d}\cdot\alpha^{j}\bar{\beta}^{d-j}$ by ($\text{by }\binom{A}{B}=\binom{A}{A-B}$
$=\binom{d}{j}\binom{j}{i+j-d}(|\mathcal{V}|-2)^{i+j-d}\cdot\alpha^{j}\bar{\beta}^{d-j}$ by ($\text{by }\binom{A}{C}\binom{A-C}{B-C}=\binom{A}{B}\binom{B}{C}$
$=\binom{d}{j}\binom{j}{d-i}(|\mathcal{V}|-2)^{i+j-d}\cdot\alpha^{j}\bar{\beta}^{d-j}$ by ($\text{by }\binom{A}{B}=\binom{A}{A-B}$
@@
An illustration of $p_{adv}-p_{A}$ graph using uniform kernel is presented in Figure 1(b).
Through the symmetrization lemma Theorem E.1, we have the following corollary, which will be used in Sec. E.3.
@@
This can be easily proved by contradiction. If the intersection part has a slope other than 1, let us assume it is $k$. Then, the slope of 1 must be either to the left or right of the axis of symmetry. Due to the symmetry, the other side must still have a slope of 1. Since the slope is a non-decreasing function of $p_{A}$, this implies that $11<k<$ 1, which leads to a contradiction. Therefore, this corollary is true. ∎
@@
Lets assume $|\mathcal{V}_{1}|\geq|\mathcal{V}_{2}|$. Denote $A_{i}(\mathcal{V})$ as the volume of $i$-th minimal trading rate. $B_{i}(\mathcal{V})$ as the corresponding volume times the trading rate. Let $r=j-i$. For the same $r$, we have the same trading rate. We calculate $\sum_{i=1}^{m}A_{i}$ by summing $v(i,i+r)$ in the order of $r$ (i.e., from larger trading rates to smaller trading rates):
@@
and for $j<d$, we have $i+j-d<i$, thus the volume
@@
$C_{\mathcal{V}}\leq\max_{C_{x},a,b}\text{ such that }\frac{(x-2)^{a}}{(x-1)^{b}}\text{ for }0\leq a<b\leq d\text{ is a monotonically decreasing function when }x>C_{x}.$
@@
Figure 1(c) illustrates the proof idea. We are using induction to prove that the blue point is always on the right side of the corresponding red point when the trading rate is less than 1.
Appendix F Tricks
This section presents some “implementation tricks” of previous defenses evaluated in this paper.

### v4

@@ D.4.3 Simple case study: dd=1,2,3
This motivate us to provide the general proof in Eq.˜11
D.5 Knapsack Solvers Yield Equivalent Results for Previous Distributions
In this section, we conduct case studies on Gaussian and Laplacian distributions, demonstrating that the results derived by knapsack solvers exactly match prior randomized smoothing results. A direct explanation is provided in Sec.˜4.2: these bounds are all black-box tight, implying their equivalence. Here, we offer an alternative perspective by deriving the results of Cohen et al. (2019) and Teng et al. (2020) using our knapsack solvers.
D.5.1 Case Study on Gaussian Distribution
For Gaussian distributions, where $p(\bm{z}|\bm{x})=\mathcal{N}(\bm{x},\sigma^{2}I)$ and $p(\bm{z}|\bm{x}_{adv})=\mathcal{N}(\bm{x}_{adv},\sigma^{2}I)$, our results are equivalent to those of Cohen et al. (2019). Following the greedy algorithm for the fractional knapsack problem (see Algorithm˜1), we select $\bm{z}$ in ascending order of the value-to-weight ratio $\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$, adding them to the set $S$ until the total weight of $S$ equals $p_{A}$, at which point the total value of items in $S$ is $p_{adv}$.
Let us define $S_{=k}=\{\bm{z}\mid\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}=k\}$ and $S_{<k}=\{\bm{z}\mid\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}<k\}$. Thus, the final result is $S=S_{<m}$ such that $\int p(\bm{z}|\bm{x})\mathbb{I}\{\bm{z}\in S_{<m}\}d\bm{z}=p_{A}$. First, observe that $S_{=k}$ forms a linear hyperplane (i.e., the boundary of $S_{<k}$ is a linear hyperplane):
$\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}=k$ $\iff\frac{\exp\left(-\frac{\|\bm{z}-\bm{x}_{adv}\|_{2}^{2}}{2\sigma^{2}}\right)}{\exp\left(-\frac{\|\bm{z}-\bm{x}\|_{2}^{2}}{2\sigma^{2}}\right)}=k$ (12)
$\iff\bm{z}^{T}(2\bm{x}_{adv}-2\bm{x})=2\sigma^{2}\log k+\|\bm{x}_{adv}\|_{2}^{2}-\|\bm{x}\|_{2}^{2}.$
This hyperplane depends on $\bm{x}_{adv}$, as its boundary is perpendicular to $\bm{x}_{adv}-\bm{x}$, indicating that the worst-case classifier depends on $\bm{x}_{adv}$. However, the final result $p_{adv}$ is determined by:
  1. 1.
    Finding $m$ such that $\int p(\bm{z}|\bm{x})\mathbb{I}\{\bm{z}\in S_{<m}\}d\bm{z}=p_{A}$. (Note that the integration result depends only on the distance between $\bm{x}$ and the hyperplane $S_{=m}$.)
  2. 2.
    Calculating $p_{adv}=\int p(\bm{z}|\bm{x}_{adv})\mathbb{I}\{\bm{z}\in S_{<m}\}d\bm{z}$. (Note that the integration result depends only on the distance between $\bm{x}_{adv}$ and the hyperplane $S_{=m}$.)
Intuive understanding of the symmetrization. To intuitively demonstrate the symmetrization across different $\bm{x}_{adv}$, we show that the distance between $S_{=k}$ and $\bm{x}$ or $\bm{x}_{adv}$ is independent of $\bm{x}_{adv}$. The distance from $S_{=k}$ to $\bm{x}$ is:
$\frac{|(2\bm{x}_{adv}-2\bm{x})^{T}\bm{x}-(2\sigma^{2}\log k+\|\bm{x}_{adv}\|_{2}^{2}-\|\bm{x}\|_{2}^{2})|}{\|2(\bm{x}_{adv}-\bm{x})\|_{2}}=\frac{|d^{2}+2\sigma^{2}\log k|}{2d},$ (13)
which is independent of $\bm{x}_{adv}$. Similarly, the distance from $S_{=k}$ to $\bm{x}_{adv}$ is:
$\frac{|(2\bm{x}_{adv}-2\bm{x})^{T}\bm{x}_{adv}-(2\sigma^{2}\log k+\|\bm{x}_{adv}\|_{2}^{2}-\|\bm{x}\|_{2}^{2})|}{\|2(\bm{x}_{adv}-\bm{x})\|_{2}}=\frac{|d^{2}-2\sigma^{2}\log k|}{2d},$ (14)
which is also independent of $\bm{x}_{adv}$. Thus, different $\bm{x}_{adv}$ yield the same $p_{adv}$, as the distances from $\bm{x}$ and $\bm{x}_{adv}$ to the hyperplane remain constant. Intuitively, as $\bm{x}_{adv}$ rotates on the sphere $\|\bm{x}_{adv}-\bm{x}\|_{2}=d$, the worst-case linear classifier rotates accordingly, but the distances from $\bm{x}$ and $\bm{x}_{adv}$ to the hyperplane remain unchanged, ensuring that the measures of the regions under $p(\bm{z}|\bm{x})$ and $p(\bm{z}|\bm{x}_{adv})$ are identical.
Deducting the result in Cohen et al. (2019). More formally, completing step 1 yields the worst-case classifier as:
$(\bm{x}_{adv}-\bm{x})^{T}\bm{z}=(\bm{x}_{adv}-\bm{x})^{T}\bm{x}+\sigma d\Phi^{-1}(p_{A}).$ (15)
Completing step 2, we obtain:
$p_{adv}=\Phi\left(\Phi^{-1}(p_{A})-\frac{d}{\sigma}\right),$ (16)
which exactly matches the result in Cohen et al. (2019) and Salman et al. (2019).
D.5.2 Case Study on Laplacian Distribution
In this section, we analyze randomized smoothing for certified robustness under L1 perturbations, assuming the noise follows a Laplacian distribution. Let the probability density functions be:
$p(\bm{z}|\bm{x})=\prod_{i=1}^{d}\frac{1}{2b}\exp\left(-\frac{|z_{i}-x_{i}|}{b}\right)=\left(\frac{1}{2b}\right)^{d}\exp\left(-\frac{\|\bm{z}-\bm{x}\|_{1}}{b}\right),$
and similarly for $p(\bm{z}|\bm{x}_{adv})$, where $\bm{x},\bm{x}_{adv}\in\mathbb{R}^{d}$ are the original and adversarial inputs, $b>0$ is the scale parameter, and $\|\cdot\|_{1}$ is the L1 norm.
Following the greedy algorithm for the fractional knapsack problem (see Algorithm˜1), we select points $\bm{z}$ in ascending order of the value-to-weight ratio $\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$, adding them to the set $S$ until the total weight of $S$ equals $p_{A}$, at which point the total value of items in $S$ is $p_{adv}$. We define:
$S_{=k}=\left\{\bm{z}\mid\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}=k\right\},\quad S_{<k}=\left\{\bm{z}\mid\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}<k\right\}.$
The goal is to find $S=S_{<m}$ such that:
$\int p(\bm{z}|\bm{x})\mathbb{I}\{\bm{z}\in S_{<m}\}d\bm{z}=p_{A},$
and then compute $p_{adv}=\int p(\bm{z}|\bm{x}_{adv})\mathbb{I}\{\bm{z}\in S_{<m}\}d\bm{z}$.
First, we compute the set $S_{=k}$:
$\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}=\frac{\exp\left(-\frac{\|\bm{z}-\bm{x}_{adv}\|_{1}}{b}\right)}{\exp\left(-\frac{\|\bm{z}-\bm{x}\|_{1}}{b}\right)}=\exp\left(\frac{\|\bm{z}-\bm{x}\|_{1}-\|\bm{z}-\bm{x}_{adv}\|_{1}}{b}\right)=k.$
Taking the natural logarithm:
$\|\bm{z}-\bm{x}\|_{1}-\|\bm{z}-\bm{x}_{adv}\|_{1}=b\log k=c.$
Thus, $S_{=k}=\{\bm{z}\mid\|\bm{z}-\bm{x}\|_{1}-\|\bm{z}-\bm{x}_{adv}\|_{1}=c\}$ is a piecewise-linear hypersurface in L1 geometry, and $S_{<k}=\{\bm{z}\mid\|\bm{z}-\bm{x}\|_{1}-\|\bm{z}-\bm{x}_{adv}\|_{1}<c\}$.
Without loss of generality, set $\bm{x}=\bm{0}$, $\bm{x}_{adv}=(d,0,\dots,0)$, where $d=\|\bm{x}_{adv}-\bm{x}\|_{1}>0$. The ratio becomes:
$\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}=\exp\left(\frac{|z_{1}|-|z_{1}-d|}{b}\right),$
since other coordinates cancel out ($|z_{i}|-|z_{i}|=0$). Define $V(z_{1})=|z_{1}|-|z_{1}-d|$. Then:
$S_{<m}=\{\bm{z}\mid V(z_{1})<c\},\quad c=b\ln m.$
Compute $V(z_{1})$:
  • If $z_{1}\leq 0$: $V(z_{1})=-z_{1}-(d-z_{1})=-d$.
  • If $0<z_{1}<d$: $V(z_{1})=z_{1}-(d-z_{1})=2z_{1}-d$.
  • If $z_{1}\geq d$: $V(z_{1})=z_{1}-(z_{1}-d)=d$.
Assuming d<c<d-d<c<d, we solve $V(z_{1})<c$. For $0<z_{1}<d$, $2z_{1}-d<c\implies z_{1}<(c+d)/2:=t$, where $t:=(c+d)/2$.
Now, compute $p_{A}=\int_{S_{<m}}p(\bm{z}|\bm{x})d\bm{z}$. Since only $z_{1}$ matters, this is the CDF of a 1D Laplacian distribution at $t$:
$p(z_{1}|x_{1}=0)=\frac{1}{2b}\exp\left(-\frac{|z_{1}|}{b}\right).$
For $t>0$:
$p_{A}=\int_{-\infty}^{t}\frac{1}{2b}\exp\left(-\frac{|z_{1}|}{b}\right)dz_{1}=\int_{-\infty}^{0}\frac{1}{2b}\exp\left(\frac{z_{1}}{b}\right)dz_{1}+\int_{0}^{t}\frac{1}{2b}\exp\left(-\frac{z_{1}}{b}\right)dz_{1}.$
Evaluate:
$\int_{-\infty}^{0}\frac{1}{2b}\exp\left(\frac{z_{1}}{b}\right)dz_{1}=\frac{1}{2},\quad\int_{0}^{t}\frac{1}{2b}\exp\left(-\frac{z_{1}}{b}\right)dz_{1}=\frac{1}{2}\left[1-\exp\left(-\frac{t}{b}\right)\right].$
Thus:
$p_{A}=\frac{1}{2}+\frac{1}{2}\left[1-\exp\left(-\frac{t}{b}\right)\right]=1-\frac{1}{2}\exp\left(-\frac{t}{b}\right).$
Solve for $t$:
$1-p_{A}=\frac{1}{2}\exp\left(-\frac{t}{b}\right)\implies\exp\left(\frac{t}{b}\right)=\frac{1}{2(1-p_{A})}\implies t=b\ln\left(\frac{1}{2(1-p_{A})}\right).$
Since $t=(c+d)/2$, we have:
$\frac{c+d}{2}=b\ln\left(\frac{1}{2(1-p_{A})}\right).$
Next, let us compute $p_{adv}=\int_{S_{<m}}p(\bm{z}|\bm{x}_{adv})d\bm{z}$, which is the CDF of Laplace($d,b$) at $t$:
$p(z_{1}|x_{adv,1}=d)=\frac{1}{2b}\exp\left(-\frac{|z_{1}-d|}{b}\right).$
We split into cases based on $t\leq d$ or $t>d$:
Case 1: $t\leq d$ (i.e., $d\geq b\ln\left(\frac{1}{2(1-p_{A})}\right)$):
$p_{adv}=\int_{-\infty}^{t}\frac{1}{2b}\exp\left(-\frac{|z_{1}-d|}{b}\right)dz_{1}=\int_{-\infty}^{t}\frac{1}{2b}\exp\left(\frac{z_{1}-d}{b}\right)dz_{1}=\frac{1}{2}\exp\left(\frac{t-d}{b}\right).$
Substitute $t=b\ln\left(\frac{1}{2(1-p_{A})}\right)$:
$p_{adv}=\frac{1}{2}\exp\left(\frac{b\ln\left(\frac{1}{2(1-p_{A})}\right)-d}{b}\right)=\frac{1}{2}\cdot\frac{1}{2(1-p_{A})}\exp\left(-\frac{d}{b}\right)=\frac{1}{4(1-p_{A})}\exp\left(-\frac{d}{b}\right).$
Case 2: $t>d$ (i.e., $d<b\ln\left(\frac{1}{2(1-p_{A})}\right)$):
$p_{adv}=\int_{-\infty}^{d}\frac{1}{2b}\exp\left(\frac{z_{1}-d}{b}\right)dz_{1}+\int_{d}^{t}\frac{1}{2b}\exp\left(-\frac{z_{1}-d}{b}\right)dz_{1}.$
Evaluate:
$\int_{-\infty}^{d}\frac{1}{2b}\exp\left(\frac{z_{1}-d}{b}\right)dz_{1}=\frac{1}{2},$
$\int_{d}^{t}\frac{1}{2b}\exp\left(-\frac{z_{1}-d}{b}\right)dz_{1}=\frac{1}{2}\left[\exp\left(-\frac{d-d}{b}\right)-\exp\left(-\frac{t-d}{b}\right)\right]=\frac{1}{2}\left[1-\exp\left(\frac{d-t}{b}\right)\right].$
So:
$p_{adv}=\frac{1}{2}+\frac{1}{2}\left[1-\exp\left(\frac{d-t}{b}\right)\right]=1-\frac{1}{2}\exp\left(\frac{d-t}{b}\right).$
Substitute $t$:
$p_{adv}=1-\frac{1}{2}\exp\left(\frac{d}{b}-\ln\left(\frac{1}{2(1-p_{A})}\right)\right)=1-\frac{1}{2}\cdot 2(1-p_{A})\exp\left(\frac{d}{b}\right)=1-(1-p_{A})\exp\left(\frac{d}{b}\right).$
Thus:
$p_{adv}=\begin{cases}1-(1-p_{A})\exp\left(\frac{d}{b}\right)&\text{if }d\leq b\ln\left(\frac{1}{2(1-p_{A})}\right),\\ \frac{1}{4(1-p_{A})}\exp\left(-\frac{d}{b}\right)&\text{otherwise}.\end{cases}$ (17)
This matches the result in Levine & Feizi (2020) and Teng et al. (2020). When $d=0$, the second case gives $p_{adv}=p_{A}$, as expected. The certified radius is obtained when $p_{adv}=0.5$, yielding $R=-b\ln(2(1-p_{A}))$.
D.6 Functional Minimization Induces Symmetrization
In this section, we provide a direct proof of why relaxing $f$ to $\mathcal{F}$ in Eq.˜2 induces symmetrization, such that solving the functional optimization $\min_{f^{\prime}\in\mathcal{F}}$ directly yields the result for input minimization. Intuitively, if a function $f^{\prime}$ performs worst on a given $\bm{x}_{adv}$, there exists another function $f^{\prime\prime}$ that performs worst on a different $\bm{x}_{adv}^{\prime}$, with both yielding equivalent results. We construct $f^{\prime\prime}$ explicitly in our proof below.
D.6.1 Case Study on Gaussian Distribution
Consider the programs:
$\min_{f^{\prime}\in\mathcal{F}}\int f^{\prime}(\bm{z})p(\bm{z}|\bm{x}_{adv})d\bm{z},\quad\text{s.t.\ }\int f^{\prime}(\bm{z})p(\bm{z}|\bm{x})d\bm{z}=p_{A},$ (18)
and
$\min_{f^{\prime}\in\mathcal{F}}\int f^{\prime}(\bm{z})p(\bm{z}|\bm{x}_{adv}^{\prime})d\bm{z},\quad\text{s.t.\ }\int f^{\prime}(\bm{z})p(\bm{z}|\bm{x})d\bm{z}=p_{A},$ (19)
where $p(\bm{z}|\bm{x})=\mathcal{N}(\bm{x},\sigma^{2}I)$, $p(\bm{z}|\bm{x}_{adv})=\mathcal{N}(\bm{x}_{adv},\sigma^{2}I)$, $p(\bm{z}|\bm{x}_{adv}^{\prime})=\mathcal{N}(\bm{x}_{adv}^{\prime},\sigma^{2}I)$, and $\|\bm{x}_{adv}-\bm{x}\|_{2}=\|\bm{x}_{adv}^{\prime}-\bm{x}\|_{2}=d$. We show that these programs yield the same result.
Without loss of generality, assume $\bm{x}=0$. There exists a rotation matrix $R$ such that $R\bm{x}_{adv}^{\prime}=\bm{x}_{adv}$ and $\det|R|=1$. For an isotropic Gaussian distribution, the density depends only on the distance to the mean, so $p(\bm{z}|0)=p(R\bm{z}|0)$ and $p(R\bm{z}|R\bm{x}_{adv}^{\prime})=p(\bm{z}|\bm{x}_{adv}^{\prime})$. Thus, Eq.˜19 is equivalent to:
$\min_{f^{\prime}\in\mathcal{F}}\int f^{\prime}(\bm{z})p(R\bm{z}|R\bm{x}_{adv}^{\prime})d\bm{z},\quad\text{s.t.\ }\int f^{\prime}(\bm{z})p(R\bm{z}|0)d\bm{z}=p_{A},$ (20)
which simplifies to:
$\min_{f^{\prime}\in\mathcal{F}}\int f^{\prime}(\bm{z})p(R\bm{z}|\bm{x}_{adv})d\bm{z},\quad\text{s.t.\ }\int f^{\prime}(\bm{z})p(R\bm{z}|0)d\bm{z}=p_{A}.$ (21)
Performing a change of variable $\bm{z}=R^{T}\bm{u}$, we obtain:
$\min_{f^{\prime}\in\mathcal{F}}\int f^{\prime}(R^{T}\bm{u})p(\bm{u}|\bm{x}_{adv})|\det R^{T}|d\bm{u},\quad\text{s.t.\ }\int f^{\prime}(R^{T}\bm{u})p(\bm{u}|0)|\det R^{T}|d\bm{u}=p_{A}.$ (22)
Since $\det|R^{T}|=1$, and defining $f^{\prime\prime}=f^{\prime}\circ R^{T}$, this becomes:
$\min_{f^{\prime\prime}\in\mathcal{F}}\int f^{\prime\prime}(\bm{u})p(\bm{u}|\bm{x}_{adv})d\bm{u},\quad\text{s.t.\ }\int f^{\prime\prime}(\bm{u})p(\bm{u}|0)d\bm{u}=p_{A},$ (23)
which is identical to Eq.˜18. Thus, the two programs yield equivalent results, confirming the symmetrization induced by relaxing $f$ to $\mathcal{F}$.
D.6.2 Case Study on Uniform Kernel
For a uniform kernel, we show that the set $S_{=k}=\{\bm{z}\mid\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}=k\}$ has the same measure under $p(\bm{z}|\bm{x})$ for all $\bm{x}_{adv}$ satisfying $\|\bm{x}_{adv}-\bm{x}\|_{0}=d$. As shown in Theorem˜5.5, the measure of $S_{=k}$ (under $p(\bm{z}|\bm{x})$) is independent of $\bm{x}_{adv}$, and thus the total value of items in $S_{=k}$ (i.e., the measure multiplied by the value-to-weight ratio) is also independent of $\bm{x}_{adv}$.
Alternatively, consider two programs:
$\min_{f^{\prime}\in\mathcal{F}}\sum_{\bm{z}}f^{\prime}(\bm{z})p(\bm{z}|\bm{x}_{adv}),\quad\text{s.t.\ }\sum_{\bm{z}}f^{\prime}(\bm{z})p(\bm{z}|\bm{x})=p_{A},$ (24)
and
$\min_{f^{\prime}\in\mathcal{F}}\sum_{\bm{z}}f^{\prime}(\bm{z})p(\bm{z}|\bm{x}_{adv}^{\prime}),\quad\text{s.t.\ }\sum_{\bm{z}}f^{\prime}(\bm{z})p(\bm{z}|\bm{x})=p_{A},$ (25)
where $\|\bm{x}_{adv}-\bm{x}\|_{0}=\|\bm{x}_{adv}^{\prime}-\bm{x}\|_{0}=d$. There exists a permutation function $P$ on token indices such that $P(\bm{x}_{adv}^{\prime})=\bm{x}_{adv}$, $P(\bm{x})=\bm{x}$, and $P$ preserves the $\ell_{0}$ distance to $\bm{x}$. For a uniform kernel, $p(\bm{z}|\bm{x})=p(P(\bm{z})|P(\bm{x}))$ for any $\bm{z}$ and $\bm{x}$, as the permutation does not map distinct tokens to the same token or identical tokens to different tokens. Thus, Eq.˜25 is equivalent to:
$\min_{f^{\prime}\in\mathcal{F}}\sum_{\bm{z}}f^{\prime}(\bm{z})p(P(\bm{z})|P(\bm{x}_{adv}^{\prime})),\quad\text{s.t.\ }\sum_{\bm{z}}f^{\prime}(\bm{z})p(P(\bm{z})|P(\bm{x}))=p_{A},$ (26)
which simplifies to:
$\min_{f^{\prime}\in\mathcal{F}}\sum_{\bm{z}}f^{\prime}(\bm{z})p(P(\bm{z})|\bm{x}_{adv}),\quad\text{s.t.\ }\sum_{\bm{z}}f^{\prime}(\bm{z})p(P(\bm{z})|\bm{x})=p_{A}.$ (27)
With a change of variable $\bm{u}=P^{-1}(\bm{z})$, this becomes:
$\min_{f^{\prime}\in\mathcal{F}}\sum_{\bm{u}}f^{\prime}(P^{-1}(\bm{u}))p(\bm{u}|\bm{x}_{adv}),\quad\text{s.t.\ }\sum_{\bm{u}}f^{\prime}(P^{-1}(\bm{u}))p(\bm{u}|\bm{x})=p_{A}.$ (28)
Defining $f^{\prime\prime}=f^{\prime}\circ P^{-1}$, this is equivalent to Eq.˜24. Thus, the two programs yield equivalent results, confirming the symmetrization for the uniform kernel.
@@ E.1 Reduction Lemma
Therefore, the certified bound of the uniform kernel is independent of the input length $N$ (dependent part only exists in network accuracy $p_{A}$), but only adversarial budget $d$. This greatly simplifies the derivation of value-to-weight ratio and volume. We give these results in the following Theorem˜5.4. We can compute the certified robustness using the uniform kernel by plugging these results into Algorithm˜1.
@@ E.2 Symmetrization Lemma
can only take $2d+1$ distinct values, specifically:
@@
$=\binom{d}{i}\binom{i}{i+j-d}(|\mathcal{V}|-2)^{i+j-d}\cdot\alpha^{j}\bar{\beta}^{d-j}$ $\text{by }\binom{A}{B}=\binom{A}{A-B}$
$=\binom{d}{i+j-d}\binom{2d-i-j}{d-j}(|\mathcal{V}|-2)^{i+j-d}\cdot\alpha^{j}\bar{\beta}^{d-j}$ $\text{by }\binom{A}{B}\binom{B}{C}=\binom{A}{C}\binom{A-C}{B-C}$
$=\binom{d}{i+j-d}\binom{2d-i-j}{d-i}(|\mathcal{V}|-2)^{i+j-d}\cdot\alpha^{j}\bar{\beta}^{d-j}$ $\text{by }\binom{A}{B}=\binom{A}{A-B}$
$=\binom{d}{j}\binom{j}{i+j-d}(|\mathcal{V}|-2)^{i+j-d}\cdot\alpha^{j}\bar{\beta}^{d-j}$ $\text{by }\binom{A}{C}\binom{A-C}{B-C}=\binom{A}{B}\binom{B}{C}$
$=\binom{d}{j}\binom{j}{d-i}(|\mathcal{V}|-2)^{i+j-d}\cdot\alpha^{j}\bar{\beta}^{d-j}$ $\text{by }\binom{A}{B}=\binom{A}{A-B}$
@@
An illustration of $p_{adv}-p_{A}$ graph using uniform kernel is presented in Figure˜1(b).
Through the symmetrization lemma Theorem˜E.1, we have the following corollary, which will be used in Sec.˜E.3.
@@
This can be easily proved by contradiction. If the intersection part has a slope other than 1, let us assume it is $k$. Then, the slope of 1 must be either to the left or right of the axis of symmetry. Due to the symmetry, the other side must still have a slope of 1. Since the slope is a non-decreasing function of $p_{A}$, this implies that $1<k<1$, which leads to a contradiction. Therefore, this corollary is true. ∎
@@
Lets assume $|\mathcal{V}_{1}|\geq|\mathcal{V}_{2}|$. Denote $A_{i}(\mathcal{V})$ as the volume of $i$-th minimal trading rate. $B_{i}(\mathcal{V})$ as the corresponding volume times the trading rate. Let r=jir=j-i. For the same $r$, we have the same trading rate. We calculate $\sum_{i=1}^{m}A_{i}$ by summing $v(i,i+r)$ in the order of $r$ (i.e., from larger trading rates to smaller trading rates):
@@
and for $j<d$, we have i+jd<ii+j-d<i, thus the volume
@@
[…]
@@
Figure˜1(c) illustrates the proof idea. We are using induction to prove that the blue point is always on the right side of the corresponding red point when the trading rate is less than 1.
Appendix F Implementation Details
This section presents some implementation tricks of previous defenses evaluated in this paper.